OpenStudy (anonymous):

(3x^3)^3= a.27x^9 b.27x^6 c.9x^9 d.9x^6

OpenStudy (anonymous):

\[(3x ^{3})^{3} = 3^{3}x ^{3*3} = 27x ^{9}\]

OpenStudy (anonymous):

why would you move the x?

OpenStudy (anonymous):

lol he didn't?

OpenStudy (anonymous):

so then why is it 3^3x?

OpenStudy (anonymous):

He used the rules of exponents. The outside ^3 is distributed to the inside 3 and 3^3 Then 3^3 = 27 x^3^3 = x^3*3 = x^9 So the final answer was 27x^9

OpenStudy (anonymous):